# Unraveling the collision scenario of the dissociative galaxy cluster Abell 56 through hydrodynamic simulations

### v2

@@ Abstract
In galaxy cluster collisions, the gas can be separated from dark matter halos. Abell 56 displays signatures of a dissociative bullet-like merger with a possible high inclination angle between the plane of orbit and the sky. Our objective is to provide a comprehensive description of the features observed in the collision scenario of Abell 56. Additionally, we aim to apply a potential weak lensing mass bias correction attributed to the merger to evaluate its impact on our findings. To investigate this, we perform tailored hydrodynamical -body simulations, varying the impact parameter. We initially identified an early scenario at Gyr after the central passage that reproduces some observational features. However, the mean temperature of keV exceeded the observed value. Our best model corresponds to the late scenario at Gyr after the pericenter, reproducing observed features of Abell 56, with an inclination of . These features include the offset of kpc between the main gas density peak and the south dark matter density peak, gas morphology, a line of sight relative velocity of km s, and a mean temperature of keV. This late model provides a plausible scenario to describe the dynamics of Abell 56. The weak lensing mass bias did not significantly impact the overall dynamics of this cluster merger.
@@
The paper is divided as follows. In Section 2, we present details about the simulation setup. In Section 3, we present the results of the collision scenario, perform analyses about the evolution of the concentration parameter, and apply the WL mass bias correction to explore the consequences in the gas morphology and mean temperature. Summary and conclusions are given in Section 4. We assume a standard CDM cosmology model with , and km s Mpc.
@@ 2 Simulation setup
To define the initial conditions, it is essential to carefully select the structural parameters of the clusters in a way that satisfies particular observational constraints. In this scenario, the A56 cluster experienced a dissociative collision, leading to potential variations in its mass over time. We chose the mass of the initial conditions in a way that ensures that the simulations match the observed data. The masses of each cluster were determined through weak-lensing analyses, resulting from the literature in M and M associated with the southern and northern galaxy subclusters respectively (Wittman et al., 2023). In the simulations, each galaxy cluster contains particles of both DM and gas.
@@
where , is the central gas density, the is the mean molecular weight per electron, is the electron number density, and is the proton mass. Assuming cm and cm, we obtained kpc and kpc.
After creating the initial conditions of the two structures with the parameters described above, we set up the collision scenario. The initial separation distance of each cluster was Mpc, with an initial approach velocity of km s. This study explored four distinct scenarios with impact parameters of kpc, kpc, kpc, and kpc.
@@ 3 Results
After numerous simulations, we obtained the so-called ‘best model’ that reproduces the morphology of A56. The best model represents the collision between the south and north substructures of A56, with an impact parameter of kpc, and km s in the initial conditions. In this section, we will discuss the collision scenarios, the influence of the impact parameter, the evolution of the concentration, and the WL mass bias.
@@ 3.1 Collision scenarios
The final temperature could, in principle, depend on the choice of certain simulation parameters. In order to attempt to decrease the temperature of the early model, we explored variations with alternative parameters. First, we considered different initial velocities, from 0 to 1200 km s. The result is that the final temperature maps (not shown) were not very sensitive to the choice of initial velocity. For the case with initial zero velocity, the temperatures are slightly smaller, but the contribution of the shock fronts remains dominant, such that the mean temperature in the relevant region is 9.4 keV. Thus, models with smaller velocities were not able to solve the issue of excessive temperature, and the model with 600 km s was adopted as the default. Secondly, we attempted to use clusters with cool cores in the initial conditions, rather than the non-cool core clusters that were assumed throughout the paper. This was achieved by setting for the gas density profile (equation 2). The steeper slope of the inner gas density profile, with the requirement of hydrostatic equilibrium, produces cool cores in the initial conditions. In the resulting cool core simulation (not shown), the central temperature was in fact somewhat decreased, but only in the region between the two DM peaks. Here again, the very hot shock fronts could not be avoided, resulting in a mean temperature of 8.2 keV within the relevant region. This result indicated that the cool core simulation could alleviate the problem, but not solve it. Thus we kept the default non-cool core models throughout the paper, and resorted to exploring the late model as the preferred candidate.
Finally, the other constraint that we aimed to satisfy in the simulation was the line of sight velocity. The line of sight velocity difference estimated by Wittman et al. (2023) between the two subclusters is km s. The early scenario only satisfies the velocity on the line of sight for near-zero angles, even when starting the simulation with a low approach velocity. This occurs due to the high relative velocity after the pericentric passage, with km s at Gyr. With such a high velocity, even a small inclination would cause the component in the line of sight to exceed the observational upper limit. In the late model, there is sufficient time after the pericentric passage to cool the gas. Additionally, with the northern cluster located only 30 kpc short of the apocenter, the relative velocity is km s at Gyr. With this much lower velocity, even the application of a large inclination angle still meets the observational constraints of the line of sight velocity. Applying an inclination angle of in the late scenario results in a line of sight velocity of km s, satisfying all the main constraints and obtaining an acceptable value of the mean temperature.
@@ 3.2 The influence of the impact parameter
Here, we will investigate the effects of different impact parameters. The impact parameter is one of the least constrained properties of the collision a priori, and it could, in principle, affect the level of dissociation. All of the collision simulations presented were conducted with an initial distance of Mpc, and the impact parameter was introduced in a perpendicular axis. For a zero impact parameter, we would have a frontal collision with a central passage. If we change the impact parameter to a nonzero value, it would induce an asymmetry in the final gas morphology. For this investigation, we conducted four different simulations with impact parameters of kpc, kpc, kpc, and kpc, as shown in Fig. 4. We initiated all simulations with a low approach velocity of km s. Consequently, the pericentric distances were kpc, kpc, kpc, and kpc, respectively. With these low pericentric distances, the variation in impact parameters did not significantly disturb the gas morphology. This occurred because the time until the pericentric passage was long enough to align the clusters with the axis of collision, which connects the centers of the clusters. Since we added the impact parameter, the time between the beginning and the first pericentric passage varies for each simulation, and it would also interfere with the impact velocity. Therefore, we needed to incline the clusters at different angles to achieve the desired distance between the DM peaks. Hence, in the second row of Fig. 4 we applied inclination angles of , , , and between the plane of the orbit and the plane of the sky, from the left to the right panel.
@@ 3.4 WL mass bias
The mass of the A56 cluster was estimated through weak lensing analysis by Wittman et al. (2023). Having undergone a recent collision, A56 is an unrelaxed cluster, and thus its NFW profile may not be well-fitted. The weak lensing analysis of Wittman et al. (2023) was based on the mass-concentration relation from Diemer & Joyce (2019). This model does not consider the previous evolutionary history of the cluster. The NFW profile and the mass-concentration relation are valid only for describing the average properties of halos (Jing, 2000). Consequently, the application of these methods to individual clusters may introduce some systematic errors, even more so in the case of merging clusters. The results of Lee et al. (2023) indicate that the difference between the weak lensing and current mass can reach as much as per cent in such collisions between M halos. This mass overestimation from the weak lensing analysis is attributed to the contraction of the halo in the merger. Consequently, there is a possible WL mass bias from the A56 merger scenario. In the following, we will discuss the effects of the WL mass bias in this hydrodynamic -body simulation.
@@
In this paper, we will not analyze the weak lensings of our simulation. Our focus is on applying the maximum WL mass bias correction results from Lee et al. (2023) in the A56 north and south clusters. This allows us to analyze the implications of this WL mass bias correction in an idealized hydrodynamic -body simulation. In previous work, Lee et al. (2023) conducted simulations to compare the virial mass and weak lensing mass over time for different mass ratio mergers. They also provided a systematic compilation of the maximum WL mass bias for each mass ratio, allowing us to consult the maximum WL mass bias for both clusters of A56. Consulting the values of the expected biases, we found that the mass for the south cluster may be overestimated by per cent and the north by per cent. We resimulated the default scenario with the WL mass bias correction to analyze the changes in our simulations. To construct the initial condition, we applied the WL mass bias correction in the for both clusters, resulting in new masses of M and M to the south and north clusters respectively. The was calculated according to Laganá et al. (2013) that estimated a baryon fraction of approximately for clusters across both mass ranges and the was obtained by the equation (6). Therefore, we set to the DM kpc and kpc, and to the gas kpc and kpc. With these parameters, we initialized the simulation with the same parameters as the default model: an initial separation of Mpc, kpc, and km s.
@@
Even applying the WL mass bias correction, the simulation results satisfied the distance constraints. In Fig. 7, the gas density morphology is presented for both cases: the original early and late scenarios in column (a), and the WL mass bias correction for each scenario in column (b). We can see that the mass peaks remain well-defined in the WL mass bias-corrected model. In the early (late) WL mass bias-corrected model, the projected distance between the DM peaks was kpc ( kpc), and the projected distance from the south DM peak to the gas peak was kpc ( kpc). To achieve this projected distance between the DM peaks in the late model, we applied an inclination angle of , between the plane of the orbit and the plane of the sky. In this scenario, the relative velocity on the line of sight was km s, still within the observational error bar of km s.
@@ 4 Conclusions
A56 is a binary galaxy cluster that underwent a dissociative merging. In our idealized simulation, we aimed to reproduce the same observational features and track the evolution of A56. Thus, we simulated distinct scenarios with different impact parameters to reach the gas morphology and the distance constraints from the observations. The projected separation between the density peaks of DM is kpc, the distance between the south cluster and the main gas density peak is kpc and the line of sight difference velocity between the two subclusters is km s (Wittman et al., 2023).
@@
We explored different collision scenarios with impact parameters of kpc, kpc, kpc, and kpc. We initialized our simulations with an initial distance of Mpc and an initial approach velocity of km s. With a low approach velocity, the pericentric distances reached were kpc, kpc, kpc, and kpc, respectively. Our default model was the one with an impact parameter of kpc because of its slightly well-defined gas density peak near the south cluster. The other models are also acceptable. For a dissociative collision, a small pericentric distance was sufficient.
@@
The late scenario, with the maximum WL-mass bias correction, achieves a well-defined gas peak at the expected position based on observational data. Despite a high inclination of between the plane of the sky and the plane of the orbit, the relative velocity on the line of sight was km s, within the observational error bar. Additionally, the new mean temperature of keV satisfies the observational constraint, close to the expected mean value of keV. Since the WL-mass bias depends on the concentration, applying this correction to the late model of A56 did not significantly alter the collision scenario. In the late model, close to the apocenter, the concentration is near its initial value. Therefore, the cluster mass does not exceed the expected range beyond the error bar because the cluster is slightly relaxed.
@@ 5 Summary
We found a late scenario, at Gyr, that successfully recovered the observational constraints of the DM peak distance and gas morphology. In this model, we identified a well-defined gas density peak near the south cluster, a mean temperature of approximately keV, and a line of sight velocity of km s, consistent with the observational constraints.

### v3

@@ Abstract
In galaxy cluster collisions, the gas can be separated from dark matter halos. Abell 56 displays signatures of a dissociative bullet-like merger with a possible high inclination angle between the plane of orbit and the sky. Our objective is to provide a comprehensive description of the features observed in the collision scenario of Abell 56. Additionally, we aim to apply a potential weak lensing mass bias correction attributed to the merger to evaluate its impact on our findings. To investigate this, we perform tailored hydrodynamical -body simulations, varying the impact parameter. We initially identified an early scenario at Gyr after the central passage that reproduces some observational features. However, the mean temperature of keV exceeded the observed value. Our best model corresponds to the late scenario at Gyr after the pericenter, reproducing observed features of Abell 56, with an inclination of . These features include the offset of kpc between the main gas density peak and the south dark matter density peak, gas morphology, a line of sight relative velocity of km s-1, and a mean temperature of keV. This late model provides a plausible scenario to describe the dynamics of Abell 56. The weak lensing mass bias did not significantly impact the overall dynamics of this cluster merger.
@@
The paper is divided as follows. In Section 2, we present details about the simulation setup. In Section 3, we present the results of the collision scenario, perform analyses about the evolution of the concentration parameter, and apply the WL mass bias correction to explore the consequences in the gas morphology and mean temperature. Summary and conclusions are given in Section 4. We assume a standard CDM cosmology model with , and km s-1 Mpc-1.
@@ 2 Simulation setup
To define the initial conditions, it is essential to carefully select the structural parameters of the clusters in a way that satisfies particular observational constraints. In this scenario, the A56 cluster experienced a dissociative collision, leading to potential variations in its mass over time. We chose the mass of the initial conditions in a way that ensures that the simulations match the observed data. The masses of each cluster were determined through weak-lensing analyses, resulting from the literature in M⊙ and M⊙ associated with the southern and northern galaxy subclusters respectively (Wittman et al., 2023). In the simulations, each galaxy cluster contains particles of both DM and gas.
@@
where , is the central gas density, the is the mean molecular weight per electron, is the electron number density, and is the proton mass. Assuming cm-3 and cm-3, we obtained kpc and kpc.
After creating the initial conditions of the two structures with the parameters described above, we set up the collision scenario. The initial separation distance of each cluster was Mpc, with an initial approach velocity of km s-1. This study explored four distinct scenarios with impact parameters of kpc, kpc, kpc, and kpc.
@@ 3 Results
After numerous simulations, we obtained the so-called ‘best model’ that reproduces the morphology of A56. The best model represents the collision between the south and north substructures of A56, with an impact parameter of kpc, and km s-1 in the initial conditions. In this section, we will discuss the collision scenarios, the influence of the impact parameter, the evolution of the concentration, and the WL mass bias.
@@ 3.1 Collision scenarios
The final temperature could, in principle, depend on the choice of certain simulation parameters. In order to attempt to decrease the temperature of the early model, we explored variations with alternative parameters. First, we considered different initial velocities, from 0 to 1200 km s-1. The result is that the final temperature maps (not shown) were not very sensitive to the choice of initial velocity. For the case with initial zero velocity, the temperatures are slightly smaller, but the contribution of the shock fronts remains dominant, such that the mean temperature in the relevant region is 9.4 keV. Thus, models with smaller velocities were not able to solve the issue of excessive temperature, and the model with 600 km s-1 was adopted as the default. Secondly, we attempted to use clusters with cool cores in the initial conditions, rather than the non-cool core clusters that were assumed throughout the paper. This was achieved by setting for the gas density profile (equation 2). The steeper slope of the inner gas density profile, with the requirement of hydrostatic equilibrium, produces cool cores in the initial conditions. In the resulting cool core simulation (not shown), the central temperature was in fact somewhat decreased, but only in the region between the two DM peaks. Here again, the very hot shock fronts could not be avoided, resulting in a mean temperature of 8.2 keV within the relevant region. This result indicated that the cool core simulation could alleviate the problem, but not solve it. Thus we kept the default non-cool core models throughout the paper, and resorted to exploring the late model as the preferred candidate.
Finally, the other constraint that we aimed to satisfy in the simulation was the line of sight velocity. The line of sight velocity difference estimated by Wittman et al. (2023) between the two subclusters is km s-1. The early scenario only satisfies the velocity on the line of sight for near-zero angles, even when starting the simulation with a low approach velocity. This occurs due to the high relative velocity after the pericentric passage, with km s-1 at Gyr. With such a high velocity, even a small inclination would cause the component in the line of sight to exceed the observational upper limit. In the late model, there is sufficient time after the pericentric passage to cool the gas. Additionally, with the northern cluster located only 30 kpc short of the apocenter, the relative velocity is km s-1 at Gyr. With this much lower velocity, even the application of a large inclination angle still meets the observational constraints of the line of sight velocity. Applying an inclination angle of in the late scenario results in a line of sight velocity of km s-1, satisfying all the main constraints and obtaining an acceptable value of the mean temperature.
@@ 3.2 The influence of the impact parameter
Here, we will investigate the effects of different impact parameters. The impact parameter is one of the least constrained properties of the collision a priori, and it could, in principle, affect the level of dissociation. All of the collision simulations presented were conducted with an initial distance of Mpc, and the impact parameter was introduced in a perpendicular axis. For a zero impact parameter, we would have a frontal collision with a central passage. If we change the impact parameter to a nonzero value, it would induce an asymmetry in the final gas morphology. For this investigation, we conducted four different simulations with impact parameters of kpc, kpc, kpc, and kpc, as shown in Fig. 4. We initiated all simulations with a low approach velocity of km s-1. Consequently, the pericentric distances were kpc, kpc, kpc, and kpc, respectively. With these low pericentric distances, the variation in impact parameters did not significantly disturb the gas morphology. This occurred because the time until the pericentric passage was long enough to align the clusters with the axis of collision, which connects the centers of the clusters. Since we added the impact parameter, the time between the beginning and the first pericentric passage varies for each simulation, and it would also interfere with the impact velocity. Therefore, we needed to incline the clusters at different angles to achieve the desired distance between the DM peaks. Hence, in the second row of Fig. 4 we applied inclination angles of , , , and between the plane of the orbit and the plane of the sky, from the left to the right panel.
@@ 3.4 WL mass bias
The mass of the A56 cluster was estimated through weak lensing analysis by Wittman et al. (2023). Having undergone a recent collision, A56 is an unrelaxed cluster, and thus its NFW profile may not be well-fitted. The weak lensing analysis of Wittman et al. (2023) was based on the mass-concentration relation from Diemer & Joyce (2019). This model does not consider the previous evolutionary history of the cluster. The NFW profile and the mass-concentration relation are valid only for describing the average properties of halos (Jing, 2000). Consequently, the application of these methods to individual clusters may introduce some systematic errors, even more so in the case of merging clusters. The results of Lee et al. (2023) indicate that the difference between the weak lensing and current mass can reach as much as per cent in such collisions between M⊙ halos. This mass overestimation from the weak lensing analysis is attributed to the contraction of the halo in the merger. Consequently, there is a possible WL mass bias from the A56 merger scenario. In the following, we will discuss the effects of the WL mass bias in this hydrodynamic -body simulation.
@@
In this paper, we will not analyze the weak lensings of our simulation. Our focus is on applying the maximum WL mass bias correction results from Lee et al. (2023) in the A56 north and south clusters. This allows us to analyze the implications of this WL mass bias correction in an idealized hydrodynamic -body simulation. In previous work, Lee et al. (2023) conducted simulations to compare the virial mass and weak lensing mass over time for different mass ratio mergers. They also provided a systematic compilation of the maximum WL mass bias for each mass ratio, allowing us to consult the maximum WL mass bias for both clusters of A56. Consulting the values of the expected biases, we found that the mass for the south cluster may be overestimated by per cent and the north by per cent. We resimulated the default scenario with the WL mass bias correction to analyze the changes in our simulations. To construct the initial condition, we applied the WL mass bias correction in the for both clusters, resulting in new masses of M⊙ and M⊙ to the south and north clusters respectively. The was calculated according to Laganá et al. (2013) that estimated a baryon fraction of approximately for clusters across both mass ranges and the was obtained by the equation (6). Therefore, we set to the DM kpc and kpc, and to the gas kpc and kpc. With these parameters, we initialized the simulation with the same parameters as the default model: an initial separation of Mpc, kpc, and km s-1.
@@
Even applying the WL mass bias correction, the simulation results satisfied the distance constraints. In Fig. 7, the gas density morphology is presented for both cases: the original early and late scenarios in column (a), and the WL mass bias correction for each scenario in column (b). We can see that the mass peaks remain well-defined in the WL mass bias-corrected model. In the early (late) WL mass bias-corrected model, the projected distance between the DM peaks was kpc ( kpc), and the projected distance from the south DM peak to the gas peak was kpc ( kpc). To achieve this projected distance between the DM peaks in the late model, we applied an inclination angle of , between the plane of the orbit and the plane of the sky. In this scenario, the relative velocity on the line of sight was km s-1, still within the observational error bar of km s-1.
@@ 4 Conclusions
A56 is a binary galaxy cluster that underwent a dissociative merging. In our idealized simulation, we aimed to reproduce the same observational features and track the evolution of A56. Thus, we simulated distinct scenarios with different impact parameters to reach the gas morphology and the distance constraints from the observations. The projected separation between the density peaks of DM is kpc, the distance between the south cluster and the main gas density peak is kpc and the line of sight difference velocity between the two subclusters is km s-1 (Wittman et al., 2023).
@@
We explored different collision scenarios with impact parameters of kpc, kpc, kpc, and kpc. We initialized our simulations with an initial distance of Mpc and an initial approach velocity of km s-1. With a low approach velocity, the pericentric distances reached were kpc, kpc, kpc, and kpc, respectively. Our default model was the one with an impact parameter of kpc because of its slightly well-defined gas density peak near the south cluster. The other models are also acceptable. For a dissociative collision, a small pericentric distance was sufficient.
@@
The late scenario, with the maximum WL-mass bias correction, achieves a well-defined gas peak at the expected position based on observational data. Despite a high inclination of between the plane of the sky and the plane of the orbit, the relative velocity on the line of sight was km s-1, within the observational error bar. Additionally, the new mean temperature of keV satisfies the observational constraint, close to the expected mean value of keV. Since the WL-mass bias depends on the concentration, applying this correction to the late model of A56 did not significantly alter the collision scenario. In the late model, close to the apocenter, the concentration is near its initial value. Therefore, the cluster mass does not exceed the expected range beyond the error bar because the cluster is slightly relaxed.
@@ 5 Summary
We found a late scenario, at Gyr, that successfully recovered the observational constraints of the DM peak distance and gas morphology. In this model, we identified a well-defined gas density peak near the south cluster, a mean temperature of approximately keV, and a line of sight velocity of km s-1, consistent with the observational constraints.